OpenStudy (anonymous):

AT A CERTAIN FACTORY, PRODUCTS ARE MADE IN VATS THAT HAVE THE CAPACITY TO HOLD 100 QUARTS. IF EACH BOTTLE OF PRODUCT CONTAINS 10/115 QUART, HOW MANY BOTTLES CAN BE MADE FROM EACH?

OpenStudy (tkhunny):

This appears to be a division problem.

OpenStudy (anonymous):

vat=100 quarts bottle=10/115 quarts answer=bottle/vats=(10/115)/100=1/1150 1150 bottles

OpenStudy (anonymous):

OKAY THANK YOU

OpenStudy (tkhunny):

Well, except that it makes no sense to do it upside down and magically come up with the final result. Converting vats to bottles: \(1\;vat\cdot\dfrac{100\;quarts}{1\;vat}\dfrac{1\;bottles}{(10/115)\;quarts} = \dfrac{100}{10/115} bottles\) One should TRY to be systematic and proceed in some logical progression of behaviors.